OpenStudy (anonymous):

The three sides of an equilateral triangle measure x + 12, 2x - 5 and 3x - 22 inches. The sides of the triangle measure

OpenStudy (anonymous):

all 3 sides of an equalateral triangle are equal. so: x+12=3x-22 34=2x x=17

OpenStudy (anonymous):

if you want to check it, then put it into all 3 equations to see if it works: 17+12=29 (2*17)-5=29 (3*17)-22=29

OpenStudy (anonymous):

Thank you. i been really struggling on that opne

OpenStudy (anonymous):

one*

OpenStudy (anonymous):

np :)